$SPAD/src/input f02aef.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
ia:=4
ib:=4
n:=4
iv:=4
a:Matrix SF:=
 [[0.5 , 1.5 , 6.6 ,  4.8 ],_
  [1.5 , 6.5 ,16.2 ,  8.6 ],_
  [6.6 ,16.2 ,37.6 ,  9.8 ],_
  [4.8 , 8.6 , 9.8 ,-17.1 ]]
b:Matrix SF:=
 [[1 , 3 , 4 , 1 ],_
  [3 ,13 ,16 ,11 ],_
  [4 ,16 ,24 ,18 ],_
  [1 ,11 ,18 ,27 ]]
result:=f02aef(ia,ib,n,iv,a,b,-1)

———-

References^{[1] nothing}

4